# Storycaster: An AI System for Immersive Room-Based Storytelling

### v1

@@ Abstract.
While Cave Automatic Virtual Environment (CAVE) systems have long enabled room-scale virtual reality and various kinds of interactivity, their content has largely remained predetermined. We present Storycaster, a generative AI CAVE system that transforms physical rooms into responsive storytelling environments. Unlike headset-based VR, Storycaster preserves spatial awareness, using live camera feeds to augment the walls with cylindrical projections, allowing users to create worlds that blend with their physical surroundings. Additionally, our system enables object-level editing, where physical items in the room can be transformed to their virtual counterparts in a story. A narrator agent guides participants, enabling them to co-create stories that evolve in response to voice commands, with each scene enhanced by generated ambient audio, dialogue, and imagery. Participants in our study ($n=13$) found the system highly immersive and engaging, with narrator and audio most impactful, while also highlighting areas for improvement in latency and image resolution.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Natural language interfacesccs: Human-centered computing User studies
@@ 1. Introduction
Spatial Augmented Reality (SAR) systems use precise geometric models of physical environments to seamlessly blend rendered virtual spaces into the real world. SAR systems such as Room Alive (Jones et al., 2014) use projection mapping techniques that remove the need for head-mounted displays (HMDs), allowing users to remain present in and directly interact with their surroundings. However, the authoring experience for these systems is difficult because the author never sees the user’s room. There is therefore a need to automatically adapt the visuals to the user’s physical environment. Recently developed AI image generation techniques seem appropriate to this task.
With the advent of large language models, we can consider the automatic generation of not just the visuals but the entire interaction itself. Furthermore, recent reliable speech recognition techniques open the possibility for the user to directly describe their desired experience in great detail, much as the characters in Star Trek: The Next Generation control and interact with the Holodeck.
Storycaster is a generative AI and SAR system designed to let users author their own experiences, augmenting physical spaces with virtual worlds. It is focused on immersive storytelling and providing users with the platform to tell any story (existing or new) and see it come to life in their own room. Storycaster distinguishes itself from virtual reality applications as its users do not lose sight of where they are and their physical surroundings. We want users to be aware of where they are, and have the ability, with a simple voice command, to change their everyday surroundings.
@@
    Room Scale Image Generation: We present an image generation pipeline that uses generative AI techniques to create projection mapping effects that alter the look of objects in the room or create entirely new scenes. The pipeline combines SDXL with Depth Control Net and a cylindrical LoRA, taking as input live depth images of the room, and a prompt to produce cylindrical images that are projected in the room. Figure 1 shows some example generations. Additionally we use Stability Audio 1.0  (Stability AI, 2024) to create matching ambient soundscapes, rendered by a Dolby spatial audio system to position audio dynamically throughout the room.
@@ 2.1. CAVE Systems
Cave Automatic Virtual Environment (CAVE) systems (Cruz-Neira et al., 1993) are projection-based systems that are designed to immerse virtual reality in a physical space. From CRT projectors to multi-projector CAVE systems with commodity hardware (Andujar et al., 2005) to the StarCAVE (DeFanti et al., 2009), they have evolved over the years, adapting through technological changes. Most recently, CAVE systems have extended beyond projectors, replacing projection screens with flat panel displays  (Febretti et al., 2013). In all cases, they are effective in having immersive room-scale VR environments that blend together virtual imagery with the user’s physical space. However, there has been limited work, to the authors’ knowledge, on CAVE systems where users can use generative AI techniques to deliver custom narratives and visuals. This is what Storycaster aims to address.
@@ 2.4. Uniting Physical and Virtual Worlds
Immersion lies on a spectrum—from augmented to virtual to mixed realities—where tangible interaction remains a central tenet (Spittle et al., 2022). Inspired by Ishii’s Tangible Bits philosophy, the goal is seamless integration between digital content and physical experience (Ishii and Ullmer, 1997). This is seen in early work with toys and augmenting these into children’s storytelling  (Glos and Umaschi, 1997). While there has been work done in scene authoring and editing in extended reality, they require users to have external screens to immerse themselves in their world  (De La Torre et al., 2024; Zhang et al., 2024; Giunchi et al., 2024; Vachha et al., 2025; Lee et al., 2025). While some storytelling systems are beginning to integrate more object-centered interactions in order to shape narratives in physical space (Balasubramani, 2024), these still lack the 360° immersion that a room provides. Work such as IllumiRoom  (Jones et al., 2013) takes a step in this direction, creating immersive experiences for gaming on TVs, turning the surroundings of the device into the environment of the game itself. Similarly, Room Alive provides a framework for turning any room into an immersive gaming experience  (Jones et al., 2014).
Storycaster is a novel step beyond these: it enables users to construct narrative worlds using simple voice commands, while spatial reasoning grounds the experience directly within physical space. In doing so, Storycaster bridges the immersive room-scale presence of CAVE systems, the narrative flexibility and co-creative potential of modern immersive storytelling, the generative power of AI-driven narrative construction, and the seamless physical and digital world integration.
By combining these threads, Storycaster seeks to redefine immersive storytelling as a physically embodied, dynamically controllable narrative experience—rooted in the user’s real environment itself.
@@ 3. Storycaster
Storycaster brings together various integrated components to create an immersive storytelling experience. The system includes: (1) a image generation pipeline that generates 360°cylindrical images from live camera feeds and maps them back on the walls and objects in the room, (2) a Narrator Agent that guides participants throughout the experience, and (3) object-editing capabilities that allow users to transform physical objects into virtual overlays contextualized within an overarching story. The overall architecture of the system is described in Figure 2, and prompts for the various modules are discussed in the Appendix.
@@ 3.1.1. Image Generation
To generate the images, a cylindrical (360°, equirectangular projection) depth image of the room is first assembled from the multiple live depth camera images. We opted for cylindrical images as this allows for a one-shot approach of generating an image that can be projected throughout the entire room, rather than generating multiple separate views. A depth ControlNet uses the cylindrical depth image as guidance to encourage image generation to follow the physical structure of the room, including the walls and furniture.
@@ 3.2. The Narrator Agent
When the Narrator prompts users for input, they can simply speak in the room. A microphone captures their voice, which is streamed via a websocket URL to Azure Speech-to-Text services for transcription. To handle pauses or hestitations without cutting off transcriptions prematurely, we increase the end-of-segmentation silence timeout to 3 seconds. Once the speech recognizer determines the user has finished speaking, it returns the final transcription, closes the stream, and plays a chime sound to signal the Narrator has successfully received the input.
@@ 3.2.3. Storyteller
Using these inputs, a detailed system prompt is crafted to generate a scene based on these personal preferences. In order to make the stories feel more like a play with heavy dialogue, we use few-shot prompting of various movie scripts from the IMDB movie database  (IMDb, [n. d.]). We hand-selected four such scripts, and extracted out the first 25 lines of two random scripts for each scene generation, ensuring variety across story structures. In particular, we aimed to emphasize setting, dialogue, and actions in the generated stories—something that movie scripts do best. Guided by these examples and detailed instructions on key storytelling principles—such as prioritizing character dialogue, using concise visual descriptions, and revealing backstory and motivations behind character’s actions—the Storyteller generates a 8-10 sentence scene.
@@ 3.2.5. Virtual Object Mapper
For object detection, we found that Grounded SAM did not consistently produce reliable masks (please see section 3.3). Instead, we opted for a more manual approach where we hand-crafted masks of various objects in the room (e.g table, ottoman, lamp, and sofas). Each mask file is labeled in the format maskobjectwhite.png, enabling straightforward retrieval.
@@ 3.2.6. Scene Audio Narrator
Additionally, background music is played in between acts through searching Spotify for top 5 playlists that meet the prompt “background instrumental music” and plays a song from the top search. This way, users do not hear the same music for each act, and in between stories.
@@ 4.3. Analysis
We used a mixed-methods approach, integrating quantitative insights from the questionnaire responses and qualitative feedback from semi-structured interviews and open-ended questionnaire items. We conducted a collaborative thematic analysis of the qualitative data, following Braun and Clarke’s (Braun and Clarke, 2006) six-phase process. Our team read and coded all interview transcripts and open-ended responses, identifying key aspects such as immersion, personalization, narrative interactivity, visual clarity, story structure, and feedback/agency. Codes were grouped into themes through iterative discussion and refinement using a shared Figma board (see Appendix 3 and supplementary materials). We defined each theme with supporting quotes and integrated them into the results, connecting qualitative insights to quantitative findings and informing our design recommendations described in Section 6.
@@ 5.1.1. Perceived Impact of Generative Aspects on Experience
In addition to these metrics, participants also provided ratings for the smoothness of the transitions between different visuals. Compared to the five dimensions, participants answers to the visual transitions were less positively received (total positive was 23.1%). The most frequent response (Mode ($\hat{x}$)), was Neutral, and the median response ($\tilde{x}$) was 3.0. The range of responses was 3.0, suggesting high variability in participant sentiment, which might have contributed to polarized “visual” dimension results.
@@ 5.1.2. Enjoyment Ratings Across Dimensions
We also analyzed enjoyment ratings to complement the dimension that enhanced the experience the most. As shown in Figure 8 and Table 2, enjoyment ratings largely mirrored the enhancement findings except for the “Object” dimension, which showed more positive trends. We subdivided this dimension in two categories 1) personal object category and 2) transformation of objects in the room. More specifically, the questions asked were: “I enjoyed seeing my object become a part of the story.” and “I enjoyed transforming objects in the environment”, which showed a more favorable distribution than “The object-level generation aspect of the room setup enhanced my experience the most.” In addition, we also asked participants to rate how much their object “played a role in the story unfolding”, which received the same total satisfaction at 62.5%.
@@ 5.2.1. Core System Feedback
Visuals: For visuals, participants said that when good, they helped place themselves in the story. However, the visuals were oftentimes vague with either unclear content, blurriness, or low-resolution images. This data is also reflected in the polarized results for the “Visual” dimension reported in the quantitative section. For instance, P5 states that when the visuals were good, “it really helped me just place myself in the story”, but that “sometimes it’s not clear, like, what is it that they’re trying to show.”. While some disliked the blurriness of some images (P11), others liked that it made it easier for them to imagine (P12). Additionally, participants had mixed reactions on the abstract nature of some of the generated images. While some participants wanted images to be more photorealistic (P9), others found the abstractness of them to be conducive to imagination. P12 compared them to comics, stating that comics are also intentionally abstract and allow people to fit the image to their own personal scenarios. A couple of participants (P10,P12) were pleasantly surprised by how artistic were the outcome visuals. Although the visuals did not match what they had initially imagined, they still found them satisfying and acceptable.
@@
Object Editing: Similar trends were seen for object-editing. While participants liked the granularity of the control and segmentation, the poor quality of the object visuals limited its overall impact. For instance, P10 said “the objects I often didn’t know what they were, but they melded it in conception”. Similarly, P10 noted how funny it was that she “said feast and it heard beast and this became covered in fur. So generally, I think I didn’t always know what the objects were supposed to be.”. Others stated that while the object-editing capability was easy to interact with, they did not feel like it added much to the story. P6 mentioned that it did not work too well but could imagine object editing being the most impactful feature.
@@
Personal Object Integration: Related to the object editing theme emerged the personal object integration. P13 enjoyed integrating a personal object into the story and found it fun and creative, “it just adds that level of personalization and customization, so I actually think that was that was very cool. I think it’s a great idea”. P7 compared it to a “MacGuffin”, referring to an object that drives the characters’ actions and serves as narrative catalyst and is a common technique in films, especially thrillers, commonly introduced in the first act. Interestingly, this coincided with the moment participants introduced their objects into the experience, although this timing was not intentionally designed. Similarly, P3 shared that their object “wasn’t a side character but felt like the main piece that kept pushing through the story“. Other participants echoed this sense of narrative integration. P6 appreciated how personal assets were brought into the story in a natural way, while P9 and P15 highlighted how the system effectively incorporated specific object-related terms (e.g., [University Name], “sea otter”) into the narrative. However, not all experiences were seamless: P15 and P11 misunderstood the input modality, expecting the system to detect objects visually rather than through voice, leading to confusion. Additionally, P14 noted that while the object was mentioned, its role felt static and lacked continuity, making it feel more programmed than immersive. These mixed reactions are reflected in the quantitative data, particularly in responses to “I enjoyed seeing my object become a part of the story.” Compared to other elements like audio or narrative, this category showed a broader spread across agreement levels.

### v2

@@ Abstract.
While Cave Automatic Virtual Environment (CAVE) systems have long enabled room-scale virtual reality and various kinds of interactivity, their content has largely remained predetermined. We present Storycaster, a generative AI CAVE system that transforms physical rooms into responsive storytelling environments. Unlike headset-based VR, Storycaster preserves spatial awareness, using live camera feeds to augment the walls with cylindrical projections, allowing users to create worlds that blend with their physical surroundings. Additionally, our system enables object-level editing, where physical items in the room can be transformed to their virtual counterparts in a story. A narrator agent guides participants, enabling them to co-create stories that evolve in response to voice commands, with each scene enhanced by generated ambient audio, dialogue, and imagery. Participants in our study ($n=13$) found the system highly immersive and engaging, identifying the narrator and audio as the most impactful elements, while also highlighting areas of improvement in latency and image resolution.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; April 13–17, 2026; Barcelona, Spainisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Natural language interfacesccs: Human-centered computing User studies
@@ 1. Introduction
Spatial Augmented Reality (SAR) systems use precise geometric models of physical environments to seamlessly blend rendered virtual spaces into the real world. Systems such as Room Alive (Jones et al., 2014) demonstrate how projection mapping techniques can transform a room without needing head-mounted displays (HMDs), allowing users to remain present in and directly interact with their surroundings. However, these kinds of experiences have traditionally relied on pre-authored content that cannot adapt dynamically to the user’s physical environment or creative goals. Authoring such experiences is difficult since designers rarely have access to the user’s room, yet convincing projection mapping requires content that is highly situated and spatially aligned with its environment.
Recently developed AI image generation techniques offer a new opportunity to address this challenge. With the advent of large language models, we can consider the automatic generation of not just the visuals but rather the entire interaction itself, from different environments, characters, and narrative structure, all responding directly to user’s creative requests. Combined with reliable speech recognition, this makes it possible for users to describe the experience they want, much as the characters in Star Trek: The Next Generation interacted with the Holodeck.
Despite advances in SAR and immersive projection, existing systems remain limited by static, pre-authored content and require specialized technical expertise to create or adapt experiences. There is a clear need for tools that empower everyday users to author dynamic, personalized stories in their own physical environments without the barriers of head-mounted displays, complex setup, or professional design skills.
Our goal with Storycaster is to explore how everyday users, such as parents, educators, and casual storytellers, might author immersive room-scale stories through natural, conversational interactions. While professional creators have traditionally authored immersive experiences, Storycaster lowers the barrier for non-experts, to enable the creation of room-scale stories through intuitive, natural interaction. We envision a future where everyday environments (living rooms, classrooms, hospital rooms), become mediums for personalized creative expression, where anyone can generate these experiences on demand without technical expertise. To clarify user roles, we distinguish between (1) Storyteller: the participant(s) who speak to the system, issue commands, and shape the narrative, (2) the AI Narrator: which guides the story, generates dialogue, and responds to user input, acting as a co-storyteller, and (3) the Audience: Anyone present in the room who experiences the story; the storyteller and AI narrator are both part of the storytelling process, and the human storyteller can also be a part of the audience.
This user-centered framing led several key design choices. First, we chose to use voice as the primary interaction modality because it is intuitive and aligns with the natural, conversational nature of storytelling. Voice input allows users to shape their narrative (and any changes they make along the way) without relying on additional devices or screens. Second, Storycaster maintains users’ awareness of their physical surroundings by augmenting the real environment, rather than replacing it. Unlike traditional virtual reality systems, which often immerse users in fully synthetic worlds, Storycaster enables users to remain present and engaged in their everyday environment, and have the ability, with a simple voice command, to change their everyday surroundings. Third, we target open-ended improvisational storytelling rather than pre-structured, choose your own adventure narratives. While the system does offer branching choices, its primary aim is not to guide users through a pre-determined narrative, but rather help them invent their own through changing settings, transforming objects, introducing characters, and redirecting the plot.
@@
    Room Scale Audio-Visual Generation: We present an image generation pipeline that uses generative AI techniques to create projection mapping effects that alter the look of objects in the room or create entirely new scenes. The pipeline combines SDXL with Depth Control Net and a cylindrical LoRA, taking as input live depth images of the room, and a prompt to produce cylindrical images that are projected in the room. Figure 1 shows some example generations. Additionally we use Stability Audio 1.0  (Stability AI, 2024) to create matching ambient soundscapes, rendered by a Dolby spatial audio system to position audio dynamically throughout the room.
@@ 2.1. CAVE Systems
Cave Automatic Virtual Environment (CAVE) systems (Cruz-Neira et al., 1993) are projection-based systems that are designed to immerse virtual reality in a physical space. From CRT projectors to multi-projector CAVE systems with commodity hardware (Andujar et al., 2005) to the StarCAVE (DeFanti et al., 2009), they have evolved over the years, adapting through technological changes. Most recently, CAVE systems have extended beyond projectors, replacing projection screens with flat panel displays  (Febretti et al., 2013). In all cases, they are effective in having immersive room-scale VR environments that blend together virtual imagery with the user’s physical space. However, there has been limited work, to the authors’ knowledge, on CAVE systems where users can use generative AI techniques to deliver custom narratives and visuals in their environments. This is what Storycaster aims to address.
@@ 2.4. Uniting Physical and Virtual Worlds
Immersion lies on a spectrum—from augmented to virtual to mixed realities—where tangible interaction remains a central tenet (Spittle et al., 2022). Inspired by Ishii’s Tangible Bits philosophy, the goal is seamless integration between digital content and physical experience (Ishii and Ullmer, 1997). This is seen in early work with toys and augmenting these into children’s storytelling  (Glos and Umaschi, 1997).
Several systems generate VR environments based on real-world physical structures. VRoamer creates VR scenes that dynamically align with unknown building layouts, enabling safe continuous walking  (Cheng et al., 2019). DreamWalker substitutes real-world walking with virtual cityscapes, overlaying virtual assets onto physical paths  (Yang et al., 2019). Sra et al.  (Sra et al., 2016) procedurally generates VR environments from reconstructed indoor spaces. All of these systems leverage physical geometry to inform environment substitution within VR, maintaining safety during movement and enabling large scale-exploration.
While Storycaster shares the idea of grounding virtual content in physical geometry, it diverges in several ways: (1) Storycaster focuses on storytelling with no headsets, where users remain visually aware of the room they are in, (2) VRoamer/DreamWalker replace the environment, while Storycaster enables users to author environments, characters, and narratives through voice commands, and (3) Storycaster includes more granular control including object-level editing to turn physical objects into narrative elements (an interaction not present in these VR systems).
Other recent AI+XR systems, such as LLMR (De La Torre et al., 2024), ImaginateAR (Lee et al., 2025), DreamCodeVR (Giunchi et al., 2024), Dreamcrafter (Vachha et al., 2025), and VRCopilot (Zhang et al., 2024), have explored scene authoring and editing in extended reality, often via natural language interfaces. However, these typically require external screens or head-mounted displays, and do not offer the in-situ, room-scale immersion that Storycaster provides. While some storytelling systems are beginning to integrate more object-centered interactions in order to shape narratives in physical space (Balasubramani, 2024), these still lack the 360° immersion that a room provides. Work such as IllumiRoom  (Jones et al., 2013) takes a step in this direction, creating immersive experiences for gaming on TVs, turning the surroundings of the device into the environment of the game itself. Similarly, Room Alive provides a framework for turning any room into an immersive gaming experience  (Jones et al., 2014).
Storycaster is a novel step beyond these: it enables users to construct narrative worlds using simple voice commands, while spatial reasoning grounds the experience directly within the physical space. In doing so, Storycaster bridges the immersive room-scale presence of CAVE systems, the narrative flexibility and co-creative potential of modern immersive storytelling, the generative power of AI-driven narrative construction, and the seamless physical and digital world integration.
By combining these threads, Storycaster seeks to redefine immersive storytelling as a physically embodied dynamically controllable narrative experience, rooted in the user’s real environment itself.
@@ 3. Storycaster
Storycaster brings together various integrated components to create an immersive storytelling experience. The system includes: (1) an image generation pipeline that generates 360° cylindrical images from live camera feeds and maps them back on the walls and objects in the room, (2) a Narrator Agent that guides participants throughout the experience, and (3) object-editing capabilities that allow users to transform physical objects into virtual overlays contextualized within an overarching story. The overall architecture of the system is described in Figure 2, and the prompts used by each module are discussed in the Appendix.
@@ 3.1. Room Features
The current implementation was installed in a single-calibrated projection-mapped room; however, the underlying approach is designed to be adaptive. Storycaster can theoretically operate in any room equipped with calibrated projectors and depth cameras, since the pipeline does not rely on room-specific characteristics or dimensions. However, environmental constraints such as cluttered spaces, or non-rectangular rooms can reduce projection quality. Future work will explore a more portable lightweight version of our setup.
@@ 3.1.1. Image Generation
To generate the images, a cylindrical (360°equirectangular projection) depth image of the room is first assembled from the multiple live depth camera images. We opted for cylindrical images as this allows for a one-shot approach of generating an image that can be projected throughout the entire room, rather than generating multiple separate views. A depth ControlNet uses the cylindrical depth image as guidance to encourage image generation to follow the physical structure of the room, including the walls and furniture.
@@
While the depth control net ensures that wall surfaces and room geometry influences image placement and perspective, the semantic meaning of physical objects is not explicitly modeled during scene-wide image synthesis, unless explicitly invoking the object-editing interaction. The prompts rather focus on describing the setting of the scene (“deep forest at dusk”, “library”) and letting the model take creative liberties with what each object becomes. This design choice allowed generations to remain robust across diverse stories but also resulted in mixed user feedback regarding object-level coherence. We discuss this further in Section  6. When users initiate object-level editing (e.g., “change the table into a black hole”), the system incorporates the semantic label of the physical object into the inpainting prompt, enabling targeted transformation. This approach allows users to personalize their stories by re-purposing familiar items as narrative anchors. Our user study found that object-level editing enhanced personalization and narrative engagement, as participants enjoyed transforming real objects into story elements (see Section  5).
@@ 3.2. The Narrator Agent
When the Narrator prompts users for input, they can simply speak in the room. A microphone captures their voice, which is streamed via a websocket URL to Azure Speech-to-Text services for transcription. To handle pauses or hesitations without cutting off transcriptions prematurely, we increase the end-of-segmentation silence timeout to 3 seconds. Once the speech recognizer determines the user has finished speaking, it returns the final transcription, closes the stream, and plays a chime sound to signal the Narrator has successfully received the input.
We recorded performance metrics to characterize system responsiveness. It took approximately 7.3 seconds to produce images (both scene-level and object-level) including upscaling, and 12 seconds for ambient audio generation. During the image generation, users were able to see the images form in the room from the initial model generation to the eventual final version after upscaling. In terms of generations, during our user study, all generations were generally successful and correspondent to user’s requests determined by their responses to an interview (see Section  5 for details). Future work includes reducing these latency values so the experience is as seamless and smooth as possible.
@@ 3.2.3. Storyteller
Using these inputs, a detailed system prompt is crafted to generate a scene based on these personal preferences. In order to make the stories feel more like a play with heavy dialogue, we use few-shot prompting of various movie scripts from the IMSDB movie script database  (IMSDb, [n. d.]). We hand-selected four such scripts from this database, and extracted out the first 25 lines of two random scripts for each scene generation, ensuring variety across story structures. In particular, we aimed to emphasize setting, dialogue, and actions in the generated stories—something that movie scripts do best. Guided by these examples and detailed instructions on key storytelling principles—such as prioritizing character dialogue, using concise visual descriptions, and revealing backstory and motivations behind character’s actions—the Storyteller generates a 8-10 sentence scene.
@@ 3.2.5. Virtual Object Mapper
For object detection, we found that Grounded SAM did not consistently produce reliable masks (please see Section 3.3). Instead, we opted for a more manual approach where we hand-crafted masks of various objects in the room (e.g., table, ottoman, lamp, and sofas). Each mask file is labeled in the format maskobjectwhite.png, enabling straightforward retrieval.
@@ 3.2.6. Scene Audio Narrator
Additionally, background music is played in between acts through searching Spotify (Spotify, 2024) for top 5 playlists that meet the prompt “background instrumental music” and plays a song from the top search. This way, users do not hear the same music for each act, and in between stories.
@@ 3.3. Object Editing
Object-level editing was included not only to demonstrate technical feasibility but to explore a new feature to support narrative making. Unlike scene-wide visual changes, object transformations allow users to re-purpose familiar physical items (sofas, tables, lamps) as narrative anchors. Even when generative quality varied, object editing created a connection between the virtual and physical worlds, enabling interactions such as turning an ottoman into a campfire, or a sofa into a boat. This form of re-contextualization is rare in prior AI + XR systems, and is a key distinction of Storycaster.
@@ 4. User Study
This evaluation was intentionally framed as a usability exploration, rather than a comprehensive benchmark against screen-based storytelling or alternate XR interfaces. Our goal was to understand how users interacted with the system, and what new forms of immersion a system like this enabled. We acknowledge the value of comparative baselines, such as 2D generative storytelling, scripted versus generative conditions, and with/without object-level editing. These comparative conditions are important for quantifying Storycaster’s unique contributions and will be the focus of future studies. In this initial work, we prioritized open-ended, in-situ interaction to capture authentic user experiences and preliminary behaviors.
@@ 4.3. Analysis
We used a mixed-methods approach, integrating quantitative insights from the questionnaire responses and qualitative feedback from semi-structured interviews and open-ended questionnaire items. We conducted a collaborative thematic analysis of the qualitative data, following Braun and Clarke’s (Braun and Clarke, 2006) six-phase process. Our team read and coded all interview transcripts and open-ended responses, identifying key aspects such as immersion, personalization, narrative interactivity, visual clarity, story structure, and feedback/agency. Codes were grouped into themes through iterative discussion and refinement using a shared Figma board (see Appendix Table 3 and Supplementary Materials). We defined each theme with supporting quotes and integrated them into the results, connecting qualitative insights to quantitative findings and informing our design recommendations described in Section 6.
@@ 5.1.1. Perceived Impact of Generative Aspects on Experience
In addition to these metrics, participants also provided ratings for the smoothness of the transitions between different visuals. Compared to the five dimensions, participants’ answers to the visual transitions were less positively received (total positive was 23.1%). The most frequent response (Mode ($\hat{x}$)), was Neutral, and the median response ($\tilde{x}$) was 3.0. The range of responses was 3.0, suggesting high variability in participant sentiment, which might have contributed to polarized “visual” dimension results.
@@ 5.1.2. Enjoyment Ratings Across Dimensions
We also analyzed enjoyment ratings to complement the dimension that enhanced the experience the most. As shown in Figure 8 and Table 2, enjoyment ratings largely mirrored the enhancement findings except for the “Object” dimension, which showed more positive trends. We subdivided this dimension in two categories: 1) personal object category and 2) transformation of objects in the room. More specifically, the questions asked were: “I enjoyed seeing my object become a part of the story.” and “I enjoyed transforming objects in the environment”, which showed a more favorable distribution than “The object-level generation aspect of the room setup enhanced my experience the most.” In addition, we also asked participants to rate how much their object “played a role in the story unfolding”, which received the same total satisfaction at 62.5%.
@@ 5.2.1. Core System Feedback
Visuals: For visuals, participants said that when good, they helped place themselves in the story. However, the visuals were oftentimes vague with either unclear content, blurriness, or low-resolution images. This data is also reflected in the polarized results for the “Visual” dimension reported in the quantitative section. For instance, P5 states that when the visuals were good, “it really helped me just place myself in the story”, but that “sometimes it’s not clear, like, what is it that they’re trying to show.”. While some disliked the blurriness of some images (P11), others liked that it made it easier for them to imagine (P12). Additionally, participants had mixed reactions on the abstract nature of some of the generated images. While some participants wanted images to be more photorealistic (P9), others found the abstractness of them to be conducive to imagination. P12 compared them to comics, stating that comics are also intentionally abstract and allow people to fit the image to their own personal scenarios. A couple of participants (P10,P12) were pleasantly surprised by how artistic the outcome visuals were. Although the visuals did not match what they had initially imagined, they still found them satisfying and acceptable.
@@
Object Editing: Similar trends were seen for object-editing. While participants liked the granularity of the control and segmentation, the poor quality of the object visuals limited its overall impact. For instance, P10 said “the objects I often didn’t know what they were, but they melded it in conception”. Similarly, P10 noted how funny it was that she “said feast and it heard beast and this became covered in fur. So generally, I think I didn’t always know what the objects were supposed to be.”. Others stated that while the object-editing capability was easy to interact with, they did not feel like it added much to the story. P6 mentioned that it did not work too well but could imagine object editing being the most impactful feature. We imagine object editing to foreground the co-presence of physical and imagined worlds, exemplifying the potential of SAR based storytelling to create new possibilities that traditional VR cannot.
@@
Personal Object Integration: Related to the object editing theme, personal object integration emerged as a key theme. P13 enjoyed integrating a personal object into the story and found it fun and creative, “it just adds that level of personalization and customization, so I actually think that was that was very cool. I think it’s a great idea”. P7 compared it to a “MacGuffin”, referring to an object that drives the characters’ actions and serves as narrative catalyst and is a common technique in films, especially thrillers, commonly introduced in the first act. Interestingly, this coincided with the moment participants introduced their objects into the experience, although this timing was not intentionally designed. Similarly, P3 shared that their object “wasn’t a side character but felt like the main piece that kept pushing through the story“. Other participants echoed this sense of narrative integration. P6 appreciated how personal assets were brought into the story in a natural way, while P9 and P15 highlighted how the system effectively incorporated specific object-related terms (e.g., [University Name], “sea otter”) into the narrative. However, not all experiences were seamless: P15 and P11 misunderstood the input modality, expecting the system to detect objects visually rather than through voice, leading to confusion. Additionally, P14 noted that while the object was mentioned, its role felt static and lacked continuity, making it feel more programmed than immersive. These mixed reactions are reflected in the quantitative data, particularly in responses to “I enjoyed seeing my object become a part of the story.” Compared to other elements like audio or narrative, this category showed a broader spread across agreement levels.